OpenStudy (anonymous):

Find the sum or difference. 5/6+1/9 A 2/5 B 8/9 C 17/18

OpenStudy (jhannybean):

\[\frac{5}{6} \cdot \frac{9}{9} + \frac{1}{9} \cdot \frac{6}{6}\]\[=\frac{45+6}{54}\]

OpenStudy (dan815):

u need common denominator to add the numerators

OpenStudy (jhannybean):

I got the LCM by finding what 9 6 equates to, that is 54. Whatever I do to the bottom of the fraction I do t the top, so I multiplied the numerators the same way.

OpenStudy (jhannybean):

Now all you need to d is simplify and reduce.

OpenStudy (jhannybean):

Good, so you get \[\frac{51}{54}\] now divide both numerator and denominator by 3, and simplify.